OpenStudy (anonymous):

Solve 7x – 2y = –3 14x + y = 14

OpenStudy (anonymous):

answer format like (4 5/7)

OpenStudy (anonymous):

x=5/7 y=4

OpenStudy (anonymous):

would it come out like (5/7,4) or (4,5/7) or what

OpenStudy (radar):

Take second equation and solve for y in terms of x. 14x + y = 14 y=14-14x Now solve by substituting for y in first equation.

OpenStudy (anonymous):

(5/7,4)

OpenStudy (radar):

7x - 2(14-14x) = -3 7x- 28 +28x = -3 35x = = -3 + 28 35x = 25 x=25/35 x=5/7

OpenStudy (radar):

From y= 14 -14x y= 14 - 14 (5/7) y = 14 -10 y=4 x=5/7 y = 4 (x,y) (5/7, 4)